# Correlated Noise in Epoch-Based Stochastic Gradient Descent: Implications for Weight Variances

### v2

@@ Conclusion
Our exploration of anti-correlations in SGD noise, which result from drawing examples without replacement, reveals a lower-than-expected weight variance in Hessian eigendirections with eigenvalues smaller than the crossover value $\lambda_{\textrm{cross}}$. The introduction of the concepts of intrinsic correlation time, shaped by the loss landscape, and constant noise correlation time, further illuminates the dynamics of SGD optimization.
One potential implication of our findings relates to the role of decreased weight variance in flat directions for SGD’s stability. This observed decrease implies a reduction in the exploration of the weight space along these directions. This diminished exploration could boost the stability of SGD, as it encourages the network to concentrate on fine-tuning parameters along a handful of directions characterized by higher curvature and weight variances.
@@ Appendix A Definition of Limiting Quantities
When we speak of a covariance matrix or an average in the main text and in the following sections of the appendix, we mean the limiting average or the limiting covariance, unless otherwise specified. In other words, we are interested in the average of a quantity over one infinite run of SGD optimization, not the mean value for a fixed update step $k$ averaged over multiple runs of SGD optimization. With this in mind, we define the covariance matrix of two quantities $\bm{\mathbf{a}}_{k}$ and $\bm{\mathbf{b}}_{k}$ as
$\textrm{cov}(\bm{\mathbf{a}}_{k},\bm{\mathbf{b}}_{k})\coloneqq\left\langle\left(\bm{\mathbf{a}}_{k}-\langle\bm{\mathbf{a}}_{k}\rangle_{k}\right)\left(\bm{\mathbf{b}}_{k}-\langle\bm{\mathbf{b}}_{k}\rangle_{k}\right)^{\top}\right\rangle_{k}$ (13)
@@
$\langle\bm{\mathbf{a}}_{k}\rangle_{k}=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}+K}\bm{\mathbf{a}}_{k}\ .$ (14)
[…]
@@
$\langle\bm{\mathbf{a}}_{k+l}\rangle$ $=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}+K}\bm{\mathbf{a}}_{k+l}$
$=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}-l}^{k_{0}-l+K}\bm{\mathbf{a}}_{k}$
$=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}-l}^{k_{0}-1}\bm{\mathbf{a}}_{k}+\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}-l+K}\bm{\mathbf{a}}_{k}\ .$ (15)
@@
$\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}-l+K}\bm{\mathbf{a}}_{k}$ $=\lim_{K\to\infty}\frac{K-l+1}{K+1}\frac{1}{K-l+1}\sum_{k=k_{0}}^{k_{0}-l+K}\bm{\mathbf{a}}_{k}$
$=\lim_{\tilde{K}\to\infty}\frac{1}{\tilde{K}+1}\sum_{k=k_{0}}^{k_{0}+\tilde{K}}\bm{\mathbf{a}}_{k}$ (16)
$\eqqcolon\langle\bm{\mathbf{a}}_{k}\rangle\ ,$ (17)
where we used $\lim_{K\to\infty}\frac{K-l+1}{K+1}=1$ and renamed $K-l$ in the second to last step. All together this gives us the desired relation $\langle\bm{\mathbf{a}}_{k}\rangle=\langle\bm{\mathbf{a}}_{k+l}\rangle$.
@@
Appendix B Weight independent Noise
It is important to note that Equation 2, the formula for the autocorrelation of the noise described in Section 4.1 and derived in Appendix E, is strictly speaking only valid for a static weight vector. During training, the weights change with each update step, which could potentially change this relationship. However, if the noise terms were independent of the weight vector, the formula would still hold. A simple case where this assumption of weight independent noise holds is a model where not only the total loss can be described by a quadratic function,
$L(\bm{\mathbf{\theta}})=\frac{1}{2}\bm{\mathbf{\theta}}^{\top}\bm{\mathbf{H}}\bm{\mathbf{\theta}}+{\rm const.}\ ,$ (18)
but the loss for a single example is described by a quadratic function as well, with the same Hessian,
$l(\bm{\mathbf{\theta}},\bm{\mathbf{x}}_{n})=\frac{1}{2}(\bm{\mathbf{\theta}}-\bm{\mathbf{\mu}}_{n})^{\top}\bm{\mathbf{H}}(\bm{\mathbf{\theta}}-\bm{\mathbf{\mu}}_{n})+{\rm const.}\ .$ (19)
However, in this model, the minimum of the loss function for this single example is offset by an example-dependent vector. This offset results in a noise term that is independent of the current weight vector for each individual example,
$\bm{\mathbf{\nabla}}\bigl{(}l(\bm{\mathbf{\theta}},\bm{\mathbf{x}}_{n})-L(\bm{\mathbf{\theta}})\bigr{)}$ $=-\bm{\mathbf{H}}\bm{\mathbf{\mu}}_{n}$
$\neq f(\bm{\mathbf{\theta}})\ .$ (20)
Therefore, also the noise introduced by SGD in this model would be independent of the current weight vector. For the later stages of training, such an assumption for the noise could be a good approximation, as evidenced by the close fit of the data to the theory in Figure 1.
Appendix C Variance Calculation
C.1 Commutativity of the covariance matrices
In this section, we show that if $\left[\bm{\mathbf{C}},\bm{\mathbf{H}}\right]=0$ also $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$ will commute with $\bm{\mathbf{C}}$, with $\bm{\mathbf{H}}$ and with each other. We make the assumptions one to three from Section 4.3 and therefore the SGD update equations become
@@ C.1 Commutativity of the covariance matrices
$\bm{\mathbf{\theta}}_{k}$ $=\left(\vec{1}-\eta\vec{H}\right)\vec{\theta}_{k-1}+\beta\vec{v}_{k-1}-\eta\vec{\delta g}_{k}\ ,$ (22)
which can be rewritten by using the vector $\vec{y}_{k}\coloneqq\bigl{(}\bm{\mathbf{\theta}}_{k}\quad\vec{v}_{k}\bigr{)}^{\top}$, combining both the current weight and velocity variable, to be
@@
$\vec{X}\coloneqq\begin{pmatrix}\vec{1}-\eta\vec{H}&\beta\vec{1}\\ -\eta\vec{H}&\beta\vec{1}\end{pmatrix}\ .$ (24)
By iteratively applying Equation 23 we obtain
@@
Under the assumption $10\leq\beta<$ 1 and $0<\eta\lambda_{i}<2(1+\beta)$, for all eigenvalues $\lambda_{i}$ of $\vec{H}$, the magnitude of the eigenvalues of $\bm{\mathbf{X}}$ will be less than one. It is straightforward to show this relation for the eigenvalues of $\vec{X}$ by using the eigenbasis of $\vec{H}$. Therefore,
@@
Because Equation 25 together with Equation 26 implies $\langle\bm{\mathbf{y}}_{k}\rangle=0$ and therefore $\langle\vec{\theta}_{k}\rangle=0$ and $\langle\vec{v}_{k}\rangle=0$, the left hand side of Equation 27 contains the covariance matrices of interest,
$\langle\vec{y}_{k}\vec{y}_{k}^{\top}\rangle=\begin{pmatrix}\vec{\Sigma}&\langle\vec{\theta}_{k}\vec{v}_{k}^{\top}\rangle\\ \langle\vec{v}_{k}\vec{\theta}_{k}^{\top}\rangle&\vec{\Sigma_{v}}\end{pmatrix}\ .$ (28)
From Equation 27 we can also infer that $\langle\vec{y}_{k}\vec{y}_{k}^{\top}\rangle$ is finite as the magnitude of the eigenvalues of $\bm{\mathbf{X}}$ is less than one. Consequently, by Equation 28, the covariance matrices $\bm{\mathbf{\Sigma}}$ and $\vec{\Sigma_{v}}$ are finite as well. The average over the noise terms $\vec{z}_{k}$ on the right hand side of Equation 27 is by assumption equal to
$\langle\vec{z}_{k+i}\vec{z}_{k+j}^{\top}\rangle=\eta^{2}\left(\delta_{i,j}-\mathbf{1}_{\{1,...,M\}}(|i-j|)\frac{M-|i-j|}{M(M-1)}\right)\cdot\begin{pmatrix}\vec{C}&\vec{C}\\ \vec{C}&\vec{C}\end{pmatrix}\ ,$ (29)
from which it follows that for any finite $h$ every matrix entry of the two by two super matrix on the right hand side of Equation 27 is a function of $\bm{\mathbf{C}}$ and $\bm{\mathbf{H}}$. Therefore, when considering the limit $h\rightarrow\infty$, $\left[\bm{\mathbf{C}},\bm{\mathbf{H}}\right]=0$ implies that $\vec{\Sigma}$ and $\vec{\Sigma_{v}}$ will also commute with $\bm{\mathbf{C}}$, with $\bm{\mathbf{H}}$ and with each other.
C.2 Proof of the variance formula for one specific eigenvalue
Since $\bm{\mathbf{\Sigma}}$ and $\bm{\mathbf{\Sigma_{v}}}$ will commute with $\bm{\mathbf{C}}$, with $\bm{\mathbf{H}}$ and with each other, it is sufficient to prove the one dimensional case. For the multidimensional case simply apply the proof in the direction of each common eigenvector individually. The expectation values discussed below are computed with respect to the asymptotic distributions of $\theta$ and $v$, since we are only interested in the asymptotic behavior of training. We want to find $\sigma_{\theta}^{2}\coloneqq\langle\theta_{k}\theta_{k}\rangle$ and $\sigma_{v}^{2}\coloneqq\langle v_{k}v_{k}\rangle$. We assume $10\leq\beta<$ 1 and $0<\eta\lambda<2(1+\beta)$ where $\lambda$ is the hessian eigenvalue.
@@ C.2 Proof of the variance formula for one specific eigenvalue
$\delta g_{k}(\theta)$ $=\delta g_{k}\textrm{ }\dots\textrm{ is independent of }\theta$ (34)
$\Rightarrow\langle\delta g_{k}\delta g_{k+h}\rangle$ $=\sigma_{\delta g}^{2}\left(\delta_{h,0}-\mathbf{1}_{\{1,...,M\}}(|h|)\frac{M-|h|}{M(M-1)}\right)$ (35)
@@
$\bm{\mathbf{x}}_{k}$ $=\vec{D}\vec{x}_{k-1}-\eta\delta g_{k}\vec{e}_{1}$ (38)
$\bm{\mathbf{x}}_{k}$ $\coloneqq\begin{pmatrix}\theta_{k}\\ \theta_{k-1}\end{pmatrix}$ (39)
$\bm{\mathbf{e}}_{1}$ $\coloneqq\begin{pmatrix}1\\ 0\end{pmatrix}$ (40)
$\vec{D}$ $\coloneqq\begin{pmatrix}1+\beta-\eta\lambda&-\beta\\ 1&0\end{pmatrix}\ .$ (41)
@@
where the second equality is due to the fact that $\langle\theta_{k}\theta_{k}\rangle=\langle\theta_{k-1}\theta_{k-1}\rangle$. As we are interested in the asymptotic covariance, this expectation value is independent of any finite shift of the index $k$. By inserting Equation 38 into $\left\langle\vec{x}_{k}\vec{x}_{k}^{\top}\right\rangle$ we arrive at the following equality
$\left\langle\vec{x}_{k}\vec{x}_{k}^{\top}\right\rangle=\vec{D}\left\langle\vec{x}_{k-1}\vec{x}_{k-1}^{\top}\right\rangle\vec{D}^{\top}+\eta^{2}\left\langle\delta g_{k}\delta g_{k}\right\rangle\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)$ (43)
@@
$\vec{\tilde{\Sigma}}-\vec{D}\vec{\tilde{\Sigma}}\vec{D}^{\top}=\eta^{2}\sigma_{\delta g}^{2}\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)\ .$ (44)
@@
$\vec{F}_{1}^{-1}$ $\coloneqq\begin{pmatrix}\eta\lambda(2-\eta\lambda)-2\beta(1+\beta-\eta\lambda)&2\beta(1+\beta-\eta\lambda)\\ -(1+\beta-\eta\lambda)&1+\beta\end{pmatrix}\ .$ (46)
@@
$\vec{F}_{2}$ $\coloneqq\begin{pmatrix}1&0\\ 2&-2\end{pmatrix}\ .$ (49)
We define the matrix $\vec{F}\coloneqq\vec{F}_{2}\vec{F}_{1}$. By applying both sides of Equation 44 to the vector $\vec{e_{1}}$, then multiplying by the matrix $\vec{F}$ from the left and using Equations 45 and 48 we obtain
$\begin{pmatrix}\sigma_{\theta}^{2}\\ \sigma_{v}^{2}\end{pmatrix}=\vec{F}\left[\eta^{2}\sigma_{\delta g}^{2}\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)\right]\vec{e}_{1}\ .$ (50)
@@
$\bm{\mathbf{F}}=\frac{1}{(1-\beta)\bigl{(}2(1+\beta)-\eta\lambda)\bigr{)}}\begin{pmatrix}\frac{1+\beta}{\eta\lambda}&\frac{2\beta(\eta\lambda-1-\beta)}{\eta\lambda}\\ 2&2(\eta\lambda-2)\end{pmatrix}\ .$ (51)
To simplify Equation 50 further we go back to Equation 38 and iterate it to obtain
$\bm{\mathbf{x}}_{k}=\vec{D}^{n}\bm{\mathbf{x}}_{k-n}-\eta\sum_{h=0}^{n-1}\vec{D}^{h}\vec{e}_{1}\delta g_{k-h}\ .$ (52)
We note that $\left\langle\bm{\mathbf{x}}_{k-n}\delta g_{k}\right\rangle=0$ for $n\geq M$. The correlation between noise terms separated by at least one epoch vanishes, and $\bm{\mathbf{x}}_{k}$ only depends on past noise terms. By setting $n=M$ we find
$\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle$ $=\vec{D}^{M}\left\langle\bm{\mathbf{x}}_{k-1-M}\delta g_{k}\right\rangle-\eta\sum_{h=0}^{M-1}\vec{D}^{h}\vec{e}_{1}\left\langle\delta g_{k}\delta g_{k-1-h}\right\rangle$
@@
where the assumption about the correlation of the noise terms, Equation 35, was inserted for the last line. Equation 53 is a sum of a finite geometric series and a derivative of that which can be simplified to
$\left\langle\bm{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle$ $=\eta\sigma_{\delta g}^{2}\frac{\bm{\mathbf{D}}^{M}+(\vec{1}-\vec{D})M-\vec{1}}{(\vec{1}-\vec{D})^{2}M(M-1)}\,\vec{e}_{1}\ .$ (54)
Substituting this result back into Equation 50 yields
$\begin{pmatrix}\sigma_{\theta}^{2}\\ \sigma_{v}^{2}\end{pmatrix}$ $=\eta^{2}\sigma_{\delta g}^{2}\vec{F}\left[\vec{e}_{1}-\left(\bm{\mathbf{E}}+\vec{E}^{\top}\right)\vec{e}_{1}\right]$ (55)
@@
$\vec{E}$ $\coloneqq\vec{D}\frac{\bm{\mathbf{D}}^{M}+(\vec{1}-\vec{D})M-\vec{1}}{(\vec{1}-\vec{D})^{2}M(M-1)}\,\vec{e}_{1}\vec{e}_{1}^{\top}\ .$ (56)
With Equation 55 we have arrived at the exact formula for the variances which can easily be evaluated numerically.
C.3 Approximation of the exact formula
It is possible to approximate the exact result for the variance assuming small or large eigenvalues, respectively. For that, it is necessary to approximate $\bm{\mathbf{D}}^{M}\bm{\mathbf{e_{1}}}$. To do so, we will use the the following eigendecomposition of $\bm{\mathbf{D}}$

### v3

@@ Conclusion
Our investigation of anti-correlations in SGD noise – arising from drawing examples without replacement – reveals a lower-than-expected weight variance in Hessian eigendirections with eigenvalues smaller than the crossover value $\lambda_{\textrm{cross}}$. By introducing the concepts of intrinsic correlation time, shaped by the loss landscape, and a constant noise correlation time, we provide deeper insights into the dynamics of SGD optimization. The reduced weight variance may allow gradients in flat directions to dominate fluctuations, steering the network toward even flatter minima. Training with anti-correlated noise can lead to improved generalization performance, suggesting that our findings may explain an important property of SGD that contributes to its success.
Note added
After completion of this work, a study by Gross et al. [41] derived a relation between weight variances and Hessian eigenvalues similar to the one discussed in this manuscript, making reference to the preprint [42] on which the present manuscript is based.
@@ Appendix A Definition of Limiting Quantities
When we speak of a covariance matrix or an average in the main text and in the following sections of the appendix, we mean the limiting average or the limiting covariance, unless otherwise specified. In other words, we are interested in the average of a quantity over one infinite run of SGD optimization, not the mean value for a fixed update step $k$ averaged over multiple runs of SGD optimization. With this in mind, we define the covariance matrix of two quantities $\boldsymbol{\mathbf{a}}_{k}$ and $\boldsymbol{\mathbf{b}}_{k}$ as
$\textrm{cov}(\boldsymbol{\mathbf{a}}_{k},\boldsymbol{\mathbf{b}}_{k})\coloneqq\left\langle\left(\boldsymbol{\mathbf{a}}_{k}-\langle\boldsymbol{\mathbf{a}}_{k}\rangle_{k}\right)\left(\boldsymbol{\mathbf{b}}_{k}-\langle\boldsymbol{\mathbf{b}}_{k}\rangle_{k}\right)^{\top}\right\rangle_{k}$ (10)
@@
$\langle\boldsymbol{\mathbf{a}}_{k}\rangle_{k}=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}+K}\boldsymbol{\mathbf{a}}_{k}\ .$ (11)
When possible, we will suppress $k$ and denote the average as $\langle\cdot\rangle$. The average is independent of the starting value $k_{0}$, therefore we can shift indices within the average, meaning $\langle\boldsymbol{\mathbf{a}}_{k}\rangle=\langle\boldsymbol{\mathbf{a}}_{k+l}\rangle$ for any $l\in\mathbb{Z}$.
@@
$\langle\boldsymbol{\mathbf{a}}_{k+l}\rangle$ $=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}+K}\boldsymbol{\mathbf{a}}_{k+l}$
$=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}-l}^{k_{0}-l+K}\boldsymbol{\mathbf{a}}_{k}$
$=\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}-l}^{k_{0}-1}\boldsymbol{\mathbf{a}}_{k}+\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}-l+K}\boldsymbol{\mathbf{a}}_{k}\ .$ (12)
@@
$\lim_{K\to\infty}\frac{1}{K+1}\sum_{k=k_{0}}^{k_{0}-l+K}\boldsymbol{\mathbf{a}}_{k}$ $=\lim_{K\to\infty}\frac{K-l+1}{K+1}\frac{1}{K-l+1}\sum_{k=k_{0}}^{k_{0}-l+K}\boldsymbol{\mathbf{a}}_{k}$
$=\lim_{\tilde{K}\to\infty}\frac{1}{\tilde{K}+1}\sum_{k=k_{0}}^{k_{0}+\tilde{K}}\boldsymbol{\mathbf{a}}_{k}$ (13)
$\eqqcolon\langle\boldsymbol{\mathbf{a}}_{k}\rangle\ ,$ (14)
where we used $\lim_{K\to\infty}\frac{K-l+1}{K+1}=1$ and renamed KlK-l in the second to last step. All together this gives us the desired relation $\langle\boldsymbol{\mathbf{a}}_{k}\rangle=\langle\boldsymbol{\mathbf{a}}_{k+l}\rangle$.
@@
Appendix B Variance Calculation
B.1 Commutativity of the covariance matrices
In this section, we show that if $\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}=\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}$ also $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$ will commute with $\boldsymbol{\mathbf{C}}$, with $\boldsymbol{\mathbf{H}}$ and with each other. We make the assumptions one to three from Section˜4.3 and therefore the SGD update equations become
@@ B.1 Commutativity of the covariance matrices
$\boldsymbol{\mathbf{\theta}}_{k}$ $=\left(\vec{1}-\eta\vec{H}\right)\vec{\theta}_{k-1}+\beta\vec{v}_{k-1}-\eta\vec{\delta g}_{k}\ ,$ (16)
which can be rewritten by using the vector $\vec{y}_{k}\coloneqq\bigl(\boldsymbol{\mathbf{\theta}}_{k}\quad\vec{v}_{k}\bigr)^{\top}$, combining both the current weight and velocity variable, to be
@@
$\vec{X}\coloneqq\begin{pmatrix}\vec{1}-\eta\vec{H}&\beta\vec{1}\\ -\eta\vec{H}&\beta\vec{1}\end{pmatrix}\$. (18)
By iteratively applying Equation˜17 we obtain
@@
Under the assumption $0\leq\beta<1$ and $0<\eta\lambda_{i}<2(1+\beta)$, for all eigenvalues $\lambda_{i}$ of $\vec{H}$, the magnitude of the eigenvalues of $\boldsymbol{\mathbf{X}}$ will be less than one. It is straightforward to show this relation for the eigenvalues of $\vec{X}$ by using the eigenbasis of $\vec{H}$. Therefore,
@@
Because Equation˜19 together with Equation˜20 implies $\langle\boldsymbol{\mathbf{y}}_{k}\rangle=0$ and therefore $\langle\vec{\theta}_{k}\rangle=0$ and $\langle\vec{v}_{k}\rangle=0$, the left hand side of Equation˜21 contains the covariance matrices of interest,
$\langle\vec{y}_{k}\vec{y}_{k}^{\top}\rangle=\begin{pmatrix}\vec{\Sigma}&\langle\vec{\theta}_{k}\vec{v}_{k}^{\top}\rangle\\ \langle\vec{v}_{k}\vec{\theta}_{k}^{\top}\rangle&\vec{\Sigma_{v}}\end{pmatrix}\$. (22)
From Equation˜21 we can also infer that $\langle\vec{y}_{k}\vec{y}_{k}^{\top}\rangle$ is finite as the magnitude of the eigenvalues of $\boldsymbol{\mathbf{X}}$ is less than one. Consequently, by Equation˜22, the covariance matrices $\boldsymbol{\mathbf{\Sigma}}$ and $\vec{\Sigma_{v}}$ are finite as well. The average over the noise terms $\vec{z}_{k}$ on the right hand side of Equation˜21 is by assumption equal to
$\langle\vec{z}_{k+i}\vec{z}_{k+j}^{\top}\rangle=\eta^{2}\left(\delta_{i,j}-\mathbf{1}_{\{1,...,M\}}(|i-j|)\frac{M-|i-j|}{M(M-1)}\right)\cdot\begin{pmatrix}\vec{C}&\vec{C}\\ \vec{C}&\vec{C}\end{pmatrix}\$, (23)
from which it follows that for any finite $h$ every matrix entry of the two by two super matrix on the right hand side of Equation˜21 is a function of $\boldsymbol{\mathbf{C}}$ and $\boldsymbol{\mathbf{H}}$. Therefore, when considering the limit $h\rightarrow\infty$, $\boldsymbol{\mathbf{C}}\boldsymbol{\mathbf{H}}=\boldsymbol{\mathbf{H}}\boldsymbol{\mathbf{C}}$ implies that $\vec{\Sigma}$ and $\vec{\Sigma_{v}}$ will also commute with $\boldsymbol{\mathbf{C}}$, with $\boldsymbol{\mathbf{H}}$ and with each other.
B.2 Proof of the variance formula for one specific eigenvalue
Since $\boldsymbol{\mathbf{\Sigma}}$ and $\boldsymbol{\mathbf{\Sigma_{v}}}$ will commute with $\boldsymbol{\mathbf{C}}$, with $\boldsymbol{\mathbf{H}}$ and with each other, it is sufficient to prove the one dimensional case. For the multidimensional case simply apply the proof in the direction of each common eigenvector individually. The expectation values discussed below are computed with respect to the asymptotic distributions of $\theta$ and $v$, since we are only interested in the asymptotic behavior of training. We want to find $\sigma_{\theta}^{2}\coloneqq\langle\theta_{k}\theta_{k}\rangle$ and $\sigma_{v}^{2}\coloneqq\langle v_{k}v_{k}\rangle$. We assume $0\leq\beta<1$ and $0<\eta\lambda<2(1+\beta)$ where $\lambda$ is the Hessian eigenvalue.
@@ B.2 Proof of the variance formula for one specific eigenvalue
$\delta g_{k}(\theta)$ $=\delta g_{k}$ (28)
$\langle\delta g_{k}\delta g_{k+h}\rangle$ $=\sigma_{\delta g}^{2}\left(\delta_{h,0}-\mathbf{1}_{\{1,...,M\}}(|h|)\frac{M-|h|}{M(M-1)}\right)$ (29)
@@
$\boldsymbol{\mathbf{x}}_{k}$ $=\vec{D}\vec{x}_{k-1}-\eta\delta g_{k}\vec{e}_{1}$ (32)
$\boldsymbol{\mathbf{x}}_{k}$ $\coloneqq\begin{pmatrix}\theta_{k}\\ \theta_{k-1}\end{pmatrix}$ (33)
$\boldsymbol{\mathbf{e}}_{1}$ $\coloneqq\begin{pmatrix}1\\ 0\end{pmatrix}$ (34)
$\vec{D}$ $\coloneqq\begin{pmatrix}1+\beta-\eta\lambda&-\beta\\ 1&0\end{pmatrix}\$. (35)
@@
where the second equality is due to the fact that $\langle\theta_{k}\theta_{k}\rangle=\langle\theta_{k-1}\theta_{k-1}\rangle$. As we are interested in the asymptotic covariance, this expectation value is independent of any finite shift of the index $k$. By inserting Equation˜32 into $\left\langle\vec{x}_{k}\vec{x}_{k}^{\top}\right\rangle$ we arrive at the following equality
$\left\langle\vec{x}_{k}\vec{x}_{k}^{\top}\right\rangle=\vec{D}\left\langle\vec{x}_{k-1}\vec{x}_{k-1}^{\top}\right\rangle\vec{D}^{\top}+\eta^{2}\left\langle\delta g_{k}\delta g_{k}\right\rangle\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)$ (37)
@@
$\vec{\tilde{\Sigma}}-\vec{D}\vec{\tilde{\Sigma}}\vec{D}^{\top}=\eta^{2}\sigma_{\delta g}^{2}\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)\ .$ (38)
@@
$\vec{F}_{1}^{-1}$ $\coloneqq\begin{pmatrix}\eta\lambda(2-\eta\lambda)-2\beta(1+\beta-\eta\lambda)&2\beta(1+\beta-\eta\lambda)\\ -(1+\beta-\eta\lambda)&1+\beta\end{pmatrix}\$. (40)
@@
$\vec{F}_{2}$ $\coloneqq\begin{pmatrix}1&0\\ 2&-2\end{pmatrix}\$. (43)
We define the matrix $\vec{F}\coloneqq\vec{F}_{2}\vec{F}_{1}$. By applying both sides of Equation˜38 to the vector $\vec{e_{1}}$, then multiplying by the matrix $\vec{F}$ from the left and using Equations˜39 and 42 we obtain
$\begin{pmatrix}\sigma_{\theta}^{2}\\ \sigma_{v}^{2}\end{pmatrix}=\vec{F}\left[\eta^{2}\sigma_{\delta g}^{2}\vec{e}_{1}\vec{e}_{1}^{\top}-\eta\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}+\left(\vec{D}\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle\vec{e}_{1}^{\top}\right)^{\top}\right)\right]\vec{e}_{1}\ .$ (44)
@@
$\boldsymbol{\mathbf{F}}=\frac{1}{(1-\beta)\bigl(2(1+\beta)-\eta\lambda)\bigr)}\begin{pmatrix}\frac{1+\beta}{\eta\lambda}&\frac{2\beta(\eta\lambda-1-\beta)}{\eta\lambda}\\ 2&2(\eta\lambda-2)\end{pmatrix}\$. (45)
To simplify Equation˜44 further we go back to Equation˜32 and iterate it to obtain
$\boldsymbol{\mathbf{x}}_{k}=\vec{D}^{n}\boldsymbol{\mathbf{x}}_{k-n}-\eta\sum_{h=0}^{n-1}\vec{D}^{h}\vec{e}_{1}\delta g_{k-h}\ .$ (46)
We note that $\left\langle\boldsymbol{\mathbf{x}}_{k-n}\delta g_{k}\right\rangle=0$ for $n\geq M$. The correlation between noise terms separated by at least one epoch vanishes, and $\boldsymbol{\mathbf{x}}_{k}$ only depends on past noise terms. By setting $n=M$ we find
$\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle$ $=\vec{D}^{M}\left\langle\boldsymbol{\mathbf{x}}_{k-1-M}\delta g_{k}\right\rangle-\eta\sum_{h=0}^{M-1}\vec{D}^{h}\vec{e}_{1}\left\langle\delta g_{k}\delta g_{k-1-h}\right\rangle$
@@
where the assumption about the correlation of the noise terms, Equation˜29, was inserted for the last line. Equation˜47 is a sum of a finite geometric series and a derivative of that which can be simplified to
$\left\langle\boldsymbol{\mathbf{x}}_{k-1}\delta g_{k}\right\rangle$ $=\eta\sigma_{\delta g}^{2}\frac{\boldsymbol{\mathbf{D}}^{M}+(\vec{1}-\vec{D})M-\vec{1}}{(\vec{1}-\vec{D})^{2}M(M-1)}\,\vec{e}_{1}\ .$ (48)
Substituting this result back into Equation˜44 yields
$\begin{pmatrix}\sigma_{\theta}^{2}\\ \sigma_{v}^{2}\end{pmatrix}$ $=\eta^{2}\sigma_{\delta g}^{2}\vec{F}\left[\vec{e}_{1}-\left(\boldsymbol{\mathbf{E}}+\vec{E}^{\top}\right)\vec{e}_{1}\right]$ (49)
@@
$\vec{E}$ $\coloneqq\vec{D}\frac{\boldsymbol{\mathbf{D}}^{M}+(\vec{1}-\vec{D})M-\vec{1}}{(\vec{1}-\vec{D})^{2}M(M-1)}\,\vec{e}_{1}\vec{e}_{1}^{\top}\ .$ (50)
With Equation˜49 we have arrived at the exact formula for the variances which can easily be evaluated numerically.
B.3 Approximation of the exact formula
It is possible to approximate the exact result for the variance assuming small or large eigenvalues, respectively. For that, it is necessary to approximate $\boldsymbol{\mathbf{D}}^{M}\boldsymbol{\mathbf{e_{1}}}$. To do so, we will use the following eigendecomposition of $\boldsymbol{\mathbf{D}}$